Sam is going to university this year. He wants to pay for his registration and buy himself a laptop so he
can do his assignments and write his online exams. Sam borrowed R15 000,00 from his grandfather at
an interest rate of 9,5% per year, compounded monthly. Determine how much Sam will owe his
grandfather after five years.
1
Expert's answer
2021-06-12T00:57:56-0400

P= R15000

I = (9.5/12)%

= 19/24%

T =5*12=60

A = p(1+r/100)t

=15000(1+(19/2400)60.

= 15000(1+0.0079)60

=15000(1.0079)60

=15000(1.605)

=R 24075.14